OpenStudy (anonymous):

the pressure (P) , temperature (T) and molar volume (V) of a gas is described by the following equation" P(V-b)=RT, where R and b are constant. Evaluate at quantity

OpenStudy (anonymous):



OpenStudy (anonymous):

at constant T

OpenStudy (irishboy123):

\[P(V-b)=RT\] \[P=\frac{RT}{(V-b)}\] can you go from here?

OpenStudy (irishboy123):

its like a normal deriv but with some funny symbols so what would \(\frac{dP}{dV}\) be, if every other letter in the equation stood for a constant?

OpenStudy (anonymous):

Will they be canceled?

OpenStudy (irishboy123):

treat the P and the V as the variables, as if it were P = P(V), and treat R,T & b as constants. normal rules apply

OpenStudy (irishboy123):

or do this first \[y=\frac{a.b}{(x-c)}\] what is dy/dx, if a,b,c are constants

OpenStudy (anonymous):

y' = ab/(x-c)^2 :)

OpenStudy (irishboy123):

y' = -ab/(x-c)^2 minus sign :p so for \[P=\frac{RT}{(V-b)}\] you would typically use this kind of symbology in thermodynamics \[ \frac{\partial P}{\partial V} \Big|_T =-\frac{RT}{(V-b)^2}\] to drive home that fact that T is being held constant